OpenStudy (anonymous):

a car travels 136 miles using 7 gallons of gas . At a rate, how far can the car travel using 42 gallons of gas

OpenStudy (anonymous):

136/7 = x/7 Here is your proportion. Can you solve for x ?

OpenStudy (anonymous):

ooops....it is supposed to be 136/7 = x/42

OpenStudy (anonymous):

Create a proportion 7/136= 42/x multiply 42*136=5712 5712/7= 816 answer 816 miles

OpenStudy (anonymous):

hope this helped

OpenStudy (anonymous):

thanks